OpenStudy (anonymous):

A parallelogram has the vertices (0, 3), (3, 0), (0, -3) and (-3, 0). Determine what type of parallelogram. Find the perimeter and area.

OpenStudy (sirm3d):

it looks like a square to me.

OpenStudy (anonymous):

Yes, I believe so too. But how do I find out the perimeter and area?

OpenStudy (anonymous):

@ghazi Help please?

OpenStudy (sirm3d):

do you know the distance formula between two points?

OpenStudy (ghazi):

 what does it look like?

OpenStudy (anonymous):

A square?

OpenStudy (ghazi):

now if you use \[D= \sqrt {(3-0)^2+(0-3)^2}=3 \sqrt 2\] for all of them, yes its a square

OpenStudy (ghazi):

also you should calculate distance between diagonals if they comes out to be sqrt 2 * side length then for sure its square :)

OpenStudy (ghazi):

@shannon.janette1 clear?

OpenStudy (anonymous):

Yes, that is clear to me!(: Can you show me how I find the perimeter and area please?

OpenStudy (ghazi):

to find the area use this formula area of square = (length of the side)^2 and perimeter = length+length+length+length=4*length

OpenStudy (anonymous):

So, it would be 3squroot2^2?

OpenStudy (ghazi):

area= \[a=(3 \sqrt 2)^2=18\]

OpenStudy (anonymous):

Okay! And perimeter would be \[\sqrt[3]{2} + \sqrt[3]{2} + \sqrt[3]{2} + \sqrt[3]{2} \right?\]

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

Okay, got it! It would be 12Squroot8 but how do I ration it out?